### v1

# Importance sampling for Sobol’ indices estimation

Abstract
We propose a new importance sampling framework for the estimation and analysis of Sobol’ indices. We show that a Sobol’ index defined under a reference input distribution can be consistently estimated from samples drawn from other sampling distributions by reweighting the estimator appropriately to account for the distribution change. We derive the optimal sampling distribution that minimizes the asymptotic variance and demonstrate its strong impact on estimation accuracy. Beyond variance reduction, the framework supports distributional sensitivity analysis via reverse importance sampling, enabling robust exploration of input distribution uncertainty with negligible additional computational cost.
1 Introduction
Complex computer models are now widely used across many scientific and engineering fields. Since they are often costly to evaluate, handling uncertainty and performing sensitivity analysis is particularly critical. In the context of global sensitivity analysis (GSA), where model inputs are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the output [8, 4].
To quantify this impact, a large variety of sensitivity indices has been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [21, 24], others are based on divergence of measures between probability distributions [14, 18, 26], while others use contrast-based functionals or Cramér–von Mises-type distances [22, 33]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [4], they are particularly well suited for scalar outputs, and can be generalized to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [1], Sobol’ indices measure the contribution of each input to the output variance. As such, they are variance-based and only capture second-order effects.
In most practical applications, the model is a complex black-box code whose output distribution is unknown. Analytical formulas for Sobol’ indices are thus rarely available. As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis. Over the last decades, many estimation strategies have been proposed, which can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [30, 20]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [3], Random Balance Designs (RBD) [13], the Effective Algorithm for Sensitivity Indices (EASI) [17], and polynomial chaos expansions [15]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [29, 25]; and (iv) nonparametric estimators—including nearest-neighbor and kernel-based methods [11, 32, 36, 38, 35] and the rank-based estimator built on Chatterjee’s correlation coefficient [41, 38]—that require only a single input–output dataset.
The theoretical properties of methods from categories (iii) and (iv) are now well established. In particular, for first-order Sobol’ indices, consistency and asymptotic normality have been proven for kernel estimators [42, 19], the Pick–Freeze estimator [25], nearest-neighbor methods [32], and rank-based estimators [41].
Among them, the kernel estimator of [19] is efficient and achieves the asymptotic optimal variance bound for estimating the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$.However, it remains delicate to implement and computationally demanding. On the other hand, the rank-based estimator of [41] is much simpler to implement but sub-optimal and only suitable for first-order indices.
A fundamental objective in Monte-Carlo estimation is to reduce the variance of the estimator without introducing bias. One of the most widely used techniques for variance reduction is importance sampling [2] (IS), which consists of modifying the sampling distribution to obtain a more efficient estimator of the quantity of interest. The variance of the IS estimator strongly depends on the choice of the sampling distribution.
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is often the main term requiring accurate estimation. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution. We prove the existence of an optimal sampling distribution that can lead to zero asymptotic variance.
To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices. A notable exception is the contribution of [23], where the authors propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a quadrature scheme (Gauss–Hermite). Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte-Carlo simulation, thereby reducing numerical effort compared to classical procedures of Sobol’ indices estimation. However, their method does not address variance minimization explicitly. In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that allows estimation from alternative sampling distributions, and also offers computational benefits presented in Section 4.
The paper is structured as follows, Section 2 introduces Sobol’ indices and the classical importance sampling framework. Section 3 presents our main theoretical results. Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real test case. Section 5 concludes and outlines future research directions.
2 Main notation
Definition of Sobol’ indices
Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ a random vector independent from $\mathbf{X}$ with known distribution. In the context of sensitivity analysis, the so-called Sobol’ index $S_{u}$ is used to quantify the impact of a subset $u\subseteq\{1,\dots,k\}$ of the input variables on the output $Y$, and is defined as:
$S_{u}=\frac{\mathrm{Var}\left(\mathbb{E}[Y\mid\mathbf{X}_{u}]\right)}{\mathrm{Var}(Y)}=\frac{\mathbb{E}\left[\mathbb{E}[Y\mid\mathbf{X}_{u}]^{2}\right]-\mathbb{E}[Y]^{2}}{\mathrm{Var}(Y)},$ (1)
where:
  • $\mathbf{X}_{u}=(X_{j})_{j\in u}$ denotes the subvector of $\mathbf{X}$ corresponding to the index set $u$,
  • $\mathbf{X}_{\bar{u}}=(X_{j})_{j\in\bar{u}}$ is the complementary subvector, i.e., the remaining components of $\mathbf{X}$.
As explained earlier, the main challenge for inference purposes lies in handling the conditional expectation $m_{u}(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}]$ in order to construct an estimator of
$\eta_{u}=\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\left[Y\cdot m_{u}(\mathbf{X}_{u})\right]$ (2)
In a given-data setting where an i.i.d. sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$ is available, optimal estimators of $\eta_{u}$ are known to be asymptotically Gaussian, with asymptotic variance achieving the efficiency bound [19] :
$\sigma^{2}_{\mathrm{opt},u}(p)=4\mathbb{E}[m_{u}^{2}(X_{u})Y^{2}]-3\mathbb{E}\left[m_{u}^{4}(X_{u})\right]-\eta_{u}^{2}.$ (3)
As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$, which drives the variance of Sobol’ index estimators. For both theoretical and practical reasons, it is often beneficial to modify the original distribution of the input variables. In the importance sampling framework, briefly introduced in the next subsection, we consider the following setting: the input vector $\mathbf{X}$ is initially distributed according to a reference distribution $p$, under which the components are assumed to be dependent, although independence can also be assumed for analytical convenience. We then consider the use of an alternative sampling distribution $q$, potentially introducing a dependency structure.Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$ and using samples $\mathbf{X}\sim q$ instead of $\mathbf{X}\sim p$.
To avoid any ambiguity, we explicitly indicate the distribution of the input $\mathbf{X}$ in the expectation notation. For example, $\mathbb{E}_{p}[\cdot]$ denotes the expectation when $\mathbf{X}$ is distributed according to the density $p$. The distribution $W$ is fixed and remains unchanged throughout. The distribution of the output $Y$ is only affected through changes in the distribution of the input.
Importance sampling (IS)
A classical variance-reduction technique. To estimate an expectation of the form $\mathbb{E}_{p}[f(\mathbf{X})]$, one rewrites it under a distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$ where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ is the likelihood ratio. To ensure the estimator is unbiased, the support of $q$ must include the support of $f(\mathbf{x})\,p(\mathbf{x})$. Given an i.i.d sample $\mathbf{X}^{(1)},...,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
$\hat{I}^{\mathrm{IS}}=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{X}^{(i)})\,w(\mathbf{X}^{(i)}).$ (4)
This estimator achieves zero variance if and only if $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost-everywhere. This ideal distribution is not usable in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as non-parametric methods [7] or parametric approaches like the cross-entropy method [34, 12].
It is worth noting that the effectiveness of IS [9] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values. The next section presents our main contributions by extending the classical IS framework to characterize the optimal sampling distribution $q$ for estimating $\eta_{u}$ and improving Sobol’ indices estimation.
3 Main results
3.1 Input distribution change via importance sampling
We consider a modification of the input distribution, applied exclusively to the random vector $\mathbf{X}\in\mathbb{R}^{k}$. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$, for any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\textbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x})$ its marginal density and introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}),$ (5)
where $q_{u}(\mathbf{x}_{u})$ and $q_{\bar{u}}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional densities, respectively. The corresponding importance weights are then given by:
$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\cdot p_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\cdot w_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}).$ (6)
Lemma
Let $q=q_{u}\cdot q_{\bar{u}}$ be any importance sampling distribution on the input $\mathbf{X}$, define
$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot Y.$ (7)
Then:
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}).$ (8)
In particular, the following identity holds:
$\eta_{u}=\mathbb{E}_{p}\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{2}\right]=\mathbb{E}_{q}\left[Z_{u}\cdot\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
Proof
See Appendix
This identity provides a reformulation of $\eta_{u}$ under the sampling distribution $q$ and can be consistently estimated using the auxiliary variable $Z_{u}$. We will show next that a suitable choice of $q$ leads to a reduced variance of the corresponding estimator.
3.2 Optimal asymptotic variance expression using importance sampling
We now investigate the case where the reference distribution of the input vector $\mathbf{X}$ can be freely modified, while the distribution of the noise variable $W$ remains fixed. We define $\phi(\mathbf{X})=\sqrt{\mathbb{E}[Y^{2}|\mathbf{X}]}$, noting that $\phi$ is model-specific and independent of the sampling distribution; its dependence on $p$ or $q$ does not need to be made explicit in the notation. Under $p$, the optimal asymptotic variance for estimating $\eta_{u}$ is given by
$\sigma^{2}_{\mathrm{opt},u}(p)=4\,\mathbb{E}_{p}\left[m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]-3\,\mathbb{E}_{p}\left[m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$ (10)
We now derive the expression of the asymptotic variance under an arbitrary auxiliary distribution $q$, and compare it to the reference case.
proposition
Let $q(\mathbf{x})$ be a sampling density over $\mathbf{X}$, and let $Z_{u}$ be defined as in Lemma 1. Then the optimal asymptotic variance is given by
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$ (11)
Using Lemma 1 and substituting the expressions of $Z_{u}$, this variance can be rewritten under $p$ as
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]$ (12)
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
This expression highlights how the asymptotic variance depends explicitly on the choice of the sampling distribution $q$ through the weights $w_{u}$ and $w_{\bar{u}}$, and through the structure of $Z_{u}$. Optimizing $q$ can therefore lead to substantial variance reduction.
3.3 Optimal distribution: sequential construction
We aim to minimize the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimizing the components of the sampling distribution $q$ in a sequential manner. Our strategy proceeds in two steps. First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$ that minimizes the variance contribution given $\mathbf{x}_{u}$. This optimal conditional density is then substituted into the variance expression. In the second step, the resulting objective is minimized with respect to the marginal $q_{u}$. To characterize the optimal distribution, we rely on a key variational lemma recalled below.
Lemma
Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$.
Then, for any $g\in\mathcal{D}_{k}$, the value $q\overset{ae}{=}g$ uniquely minimizes the functional
$q\mapsto\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x}\quad,\quad q\in\mathcal{D}_{k}.$
Proof
Let $q\in\mathcal{D}_{k}$. By Jensen’s inequality applied to the convex function $t\mapsto t^{2}$, we obtain:
[…]
Where equality is achieved if and only if $q=g$ almost everywhere.
First step: optimization of $q_{\bar{u}\mid u}$
We begin by optimizing the conditional component $q_{\bar{u}\mid u}$ of $q$, assuming that $q_{u}$ is fixed. From Proposition 3.2, the conditional distribution influences the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ in eq(3.6) only through the term
$\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
Minimizing this expectation is therefore sufficient, since it is the only component of the variance that depends on $q_{\bar{u}\mid u}$. Applying Lemma 2 conditionally on $\mathbf{X}_{u}$ gives the optimal conditional density:
$q^{*}_{\bar{u}\mid u}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})=\frac{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi(\mathbf{X})}{\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}\right]},$ (13)
where the denominator ensures normalization. With this choice, the variance contribution when $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$ becomes $\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{\bar{u}\mid u}\sim p_{\bar{u}}$ (i.e., $w_{\bar{u}\mid u}=1$). A full derivation is provided in the Appendix. Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality confirms that
$\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
which guarantees that the optimal conditional distribution $q^{*}_{\bar{u}\mid u}$ systematically reduces this part of the variance.
Second step: optimisation of the marginal $q_{u}$
We now fix the component $q_{\bar{u}\mid u}$, and seek to minimize the asymptotic variance over the marginal $q_{u}$. For any such fixed conditional, the sampling density is of the form $q=q_{u}\cdot q_{\bar{u}\mid u}$, and the corresponding variance reads
$\sigma_{\operatorname{opt},u}^{2}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{\bar{u}\mid u}$. The optimal marginal minimizing this expression is given by
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p(\mathbf{x}_{u})\,\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\operatorname{opt},u}(q)=\left(\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
We distinguish below two particular cases:
  • Case A – Reference conditional: $q_{\bar{u}\mid u}=p_{\bar{u}\mid u}$
    $S_{A}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-3\,m_{u}^{4}(\mathbf{x}_{u}).$
  • Case B – Optimal conditional: $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$
    $S_{B}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}-3\,m_{u}^{4}(\mathbf{x}_{u}).$
This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, shown in step one. The full chain of inequalities then reads
$\sigma^{2}_{\operatorname{opt},u}(q^{*})\leq\sigma^{2}_{\operatorname{opt},u}(q_{u}^{*}\cdot p_{\bar{u}\mid u})\leq\sigma^{2}_{\operatorname{opt},u}(p),$
this follows from Jensen’s inequality:
$\left(\mathbb{E}_{p}[\sqrt{S_{B}(\mathbf{X}_{u})}]\right)^{2}\leq\left(\mathbb{E}_{p}[\sqrt{S_{A}(\mathbf{X}_{u})}]\right)^{2}\leq\mathbb{E}_{p}[S_{A}(\mathbf{X}_{u})].$
That is, joint optimisation in Case B achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
3.4 Zero-variance optimal distribution
In the idealised case where the function $f$ is fully deterministic with all inputs controllable (i.e., there is no external noise variable $W$ and $Y=f(\mathbf{X})$), it becomes theoretically possible to achieve zero-variance in the estimation of $\eta_{u}$. This setting can be interpreted as a special case of the general framework discussed earlier. We still proceed in two steps, but expressions simplify significantly. In particular, the asymptotic variance now takes the form
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y^{2}\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
Applying Lemma 2 sequentially yields:
$q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (16)
The resulting fully adapted sampling density is
$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\cdot q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m_{u}(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (17)
Hence, the density $q^{*}(\mathbf{x})$ achieves a zero-variance estimator in theory. However, this ideal result is not achievable in practice, as the optimal distribution depends on the unknown quantity $\eta_{u}$ that we aim to estimate.
4 Numerical Experiments
4.1 Analytical Illustration of Variance Reduction via IS
In this section, we illustrate the effectiveness of an IS transformation in reducing the variance $\eta_{u}$ estimators. As a benchmark example, we consider the classical Sobol’ g-function [8]:
$g(X_{1},\dots,X_{k})=\prod_{i=1}^{k}\frac{|4X_{i}-2|+a_{i}}{1+a_{i}},$ (18)
where $X_{i}\sim\mathcal{U}[0,1]$ are assumed to be independent inputs. This function is widely used in sensitivity analysis due to its non-linear, non-monotonic nature and the availability of analytical Sobol’ indices. We consider a 3-dimensional case $(k=3)$ with $a_{i}=i$, and focus on estimating the second-order Sobol’ index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$. The third input $X_{3}$ is treated as an uncontrolled variable $W$. We consider the case A in (3.3), where $\bar{u}=\emptyset$. In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically, including the optimal sampling density. Specifically, the optimal density $q^{\star}$ to estimate $\eta_{12}$ is given by
$q^{*}(x_{1},x_{2})=\frac{p(x_{1},x_{2})\cdot\sqrt{S(x_{1},x_{2})}}{\mathbb{E}_{p}[\sqrt{S(X_{1},X_{2})}]},\quad\text{with}\quad S(x_{1},x_{2})=4\,m^{2}(x_{1},x_{2})\,\phi^{2}(x_{1},x_{2})-3\,m^{4}(x_{1},x_{2}).$
The conditional expectation $m(x_{1},x_{2})=\mathbb{E}[g(x_{1},x_{2},X_{3})\mid X_{1}=x_{1},X_{2}=x_{2}]$, and the squared term $\phi^{2}(x_{1},x_{2})=\mathbb{E}[g^{2}(x_{1},x_{2},X_{3})\mid X_{1}=x_{1},X_{2}=x_{2}]$, can be derived analytically as
$m(x_{1},x_{2})=\frac{(|4x_{1}-2|+1)(|4x_{2}-2|+2)}{6},\quad\phi^{2}(x_{1},x_{2})=\frac{99}{96}\cdot m^{2}(x_{1},x_{2}).$
This leads to the simplified expression
$S(x_{1},x_{2})=\frac{9}{8}\cdot m^{4}(x_{1},x_{2}),\quad\text{so that}\quad q^{*}(x_{1},x_{2})=\frac{m^{2}(x_{1},x_{2})}{\eta_{\{1,2\}}}.$
We first plot the shape of $q^{*}$ over $[0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
$q_{t}(x_{1},x_{2})=(1-t)\,p(x_{1},x_{2})+t\,q^{*}(x_{1},x_{2}),\quad t\in[0,1].$
The coefficient of variation shown in Figure 1 is derived from the asymptotic variance under IS. It is defined as
$\mathrm{CV}_{\mathrm{opt}}(t)=\frac{\sigma_{\mathrm{opt},u}(q_{t})}{\eta_{\{1,2\}}},\quad\text{with}\quad\sigma^{2}_{\mathrm{opt},u}(q_{t})=\mathbb{E}_{p}\left[\frac{S(X_{1},X_{2})}{q_{t}(X_{1},X_{2})}\right]-\eta_{\{1,2\}}^{2}.$
Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte-Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
(a) Optimal density $q^{\star}(x_{1},x_{2})$
(b) Coefficient of variation as a function of $t$
Figure 1: Left: optimal sampling density $q^{\star}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{\star}$.
Parametric optimization
An alternative strategy consists of introducing a parametric family of auxiliary sampling densities for $(X_{1},X_{2})$, defined by
$q_{\alpha,\beta}(x_{1},x_{2})=\mathrm{Beta}(x_{1};\alpha_{1},\beta_{1})\cdot\mathrm{Beta}(x_{2};\alpha_{2},\beta_{2}),$
where the shape parameters $\alpha_{i},\beta_{i}$ are positive real numbers, i.e., $\alpha_{i},\beta_{i}\in\mathbb{R}_{+}^{*}$ for $i=1,2$. We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform distribution at $(\alpha,\beta)=(1,1)$. The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimizes the asymptotic variance, namely
$(\alpha^{*},\beta^{*})=\arg\min_{(\alpha,\beta)\in(\mathbb{R}_{+}^{*})^{2}}\sigma^{2}_{\mathrm{opt},u}(q_{\alpha,\beta}).$
The following plot shows how the asymptotic variance changes across the Beta family and highlights the optimal region.
Figure 2: Asymptotic variance $\sigma_{\mathrm{opt}}^{2}(\alpha,\beta)$ over the Beta parameter grid (symmetric case) displayed on a logarithmic scale.
[…]
4.2 Distributional Sensitivity via Reverse Importance Sampling
We study the sensitivity of the quantity
$\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m_{u}^{2}(\mathbf{X}_{u})\right],$
with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$. This falls into the framework of reverse importance sampling; this method, originally referred to as importance sampling sensitivity analysis [6], is now often described as a reverse importance sampling mechanism [27] and exploits the fact that a single set of observations can be reused to estimate responses under multiple reference distributions. Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [5].
The rank estimator [41], based on the sequential correlation of outputs, takes as input a sequence $Y^{(i)}$ sorted with respect to the associated inputs $\mathbf{X}_{u}^{(i)}$, and converges to $\eta_{u}$ via the average of consecutive products. While initially defined under the distribution $p$, this estimator can be naturally extended to estimate $\eta_{u}(\theta)$ under any distribution $p_{\theta}$, using the results from Section 3.1, the convergence to $\eta_{u}(\theta)$ is guaranteed.
For each $\theta$, we define the reweighted output
$Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\cdot\frac{p_{\theta,\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}\cdot Y,$ (19)
so that the rank-based estimator becomes
$\hat{\eta}^{\text{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)}.$ (20)
Hence, from a single dataset generated under $p$, one can efficiently explore the sensitivity of $\eta_{u}$ (and thus the associated Sobol’ index $S_{u}$) across the entire family of distributions $p_{\theta}$, without performing any additional evaluations of the black-box model.
Illustrative test case
Physical model
[…]
$w_{\boldsymbol{\theta}}(\mathbf{X})=\prod_{j=1}^{4}\frac{\operatorname{Beta}_{\theta_{j}}(X_{j}^{\mathrm{std}})}{\operatorname{Unif}[0,1](X_{j}^{\mathrm{std}})}.$ (21)
With the rank estimator in eq(4.3), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$. Two perturbation strategies are considered: (i) marginal perturbation, where only the pair $\theta_{j}=(\alpha_{j},\beta_{j})$ associated with a single input $X_{j}$ is varied while the remaining marginals remain uniform; and (ii) global perturbation, where several or all input marginals are varied simultaneously.
To keep the presentation simple, we focus on marginal perturbations. Specifically, we assess the distributional sensitivity of the contrast $\eta_{u}$ with $u=\{4\}$ by varying $\theta_{j}$ for one input variable at a time, keeping all others unchanged. This yields a family of perturbed estimates $\eta_{4}(\theta_{j})$, each revealing how the index $\eta_{4}$ responds to changes in the distribution of a specific input $X_{j}$, for $j=1,\dots,4$. All computations are based on the same reference dataset.
(a) Varying $\alpha_{1},\beta_{1}$ of $X_{1}$
(b) Varying $\alpha_{2},\beta_{2}$ of $X_{2}$
(c) Varying $\alpha_{3},\beta_{3}$ of $X_{3}$
(d) Varying $\alpha_{4},\beta_{4}$ of $X_{4}$
Figure 3: Effect of marginal distribution changes on $\eta_{4}$: each subplot shows how $\eta_{4}$ evolves as the Beta parameters of one input $X_{j}$ are varied, while others are kept uniform. All curves are computed from the same baseline sample ($N=2500$).
These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, with markedly different behaviors across the four inputs. Overall, these confirm that the proposed given-data IS framework offers a powerful and cost-effective mean to perform detailed distributional local sensitivity analyses of Sobol’ indices.
5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution different from the reference one. In particular, we show that by choosing optimal densities, it is possible to achieve a lower bound on the asymptotic variance than the efficiency bound established in [19]. In some special cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. We also show that combining our method with the rank-based estimator [41] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [27]. Our approach can be naturally extended to other contexts, such as the given-model setting where a computational budget is available to evaluate the black-box model. In such cases, we expect to approximate the optimal densities presented in our framework and achieve significant variance reduction. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
Appendix A Lemma 1 and other derivations
Proof of Lemma 1
We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{p}\left[\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}),$
and
$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m_{u}^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m_{u}^{2}(\mathbf{X}_{u})]$
Derivation of the optimal variance under $p_{u}$ in (3.7)
We start from the expression
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$
Using the last identity we deduce
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}=w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u}),\quad\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}=w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u}).$
For the second term $\mathbb{E}_{q_{\bar{u}}}[Z_{u}^{2}\mid\mathbf{X}_{u}]$, we write
$\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]$ $=\mathbb{E}_{q}\left[\left(\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y\right)^{2}\middle|\mathbf{X}_{u}\right]$
$=w_{u}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
Combining all terms, we obtain:
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
Computing optimal variance bounds
Plugging $q^{*}_{\bar{u}\mid u}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{\bar{u}\mid u}))$. We obtain:
$\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]$ $=\mathbb{E}_{p}\left[\frac{p(\mathbf{X})}{q^{*}(\mathbf{X}\mid\mathbf{X}_{u})}\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]$
$=\mathbb{E}_{p}\left[\frac{\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]}{\phi(\mathbf{X})}\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]$
$=\mathbb{E}_{p}\left[\phi(\mathbf{X})\cdot\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}.$
Similarly for $q^{*}(\mathbf{x}_{u})$:
$\mathbb{E}_{p}\left[w(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$ $=\mathbb{E}_{p}\left[\frac{p(\mathbf{X}_{u})}{q^{*}(\mathbf{X}_{u})}\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]}{\sqrt{S(\mathbf{X}_{u})}}\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\cdot\sqrt{S(\mathbf{X}_{u})}\right]-\eta_{u}^{2}$
$=\left(\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\right)^{2}-\eta_{u}^{2}.$
Derivation of the zero-variance formula
We start from the general expression:
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
Now Plugging both optimal densities simultaneously:
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}^{2}(\mathbf{x}_{u})}$
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\left(4\,m^{2}(\mathbf{X})\cdot\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}\right]-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot\left(4\,m^{2}(\mathbf{X})\cdot\mathbb{E}_{p}\left[\frac{m(\mathbf{X})}{f(\mathbf{X})}\cdot f^{2}(\mathbf{X})\mid\mathbf{X}\right]-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot\left(4\,m^{3}(\mathbf{X})\cdot m(\mathbf{X})-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot m^{4}(\mathbf{X})\right]-\eta_{u}^{2}$
$=\eta_{u}\cdot\mathbb{E}_{p}\left[m^{2}(\mathbf{X})\right]-\eta_{u}^{2}$
$=\eta_{u}^{2}-\eta_{u}^{2}=0.$

### v2

@@
 We propose a new importance sampling framework for the estimation and analysis of Sobol’ indices.
-We show that a Sobol’ index defined under a reference input distribution can be consistently estimated from samples drawn from other sampling distributions by reweighting the estimator appropriately to account for the distribution change.
-We derive the optimal sampling distribution that minimizes the asymptotic variance and demonstrate its strong impact on estimation accuracy.
-Beyond variance reduction, the framework supports distributional sensitivity analysis via reverse importance sampling, enabling robust exploration of input distribution uncertainty with negligible additional computational cost.
+We focus on the estimation of the conditional second-moment quantity underlying these indices, which is the most challenging term to estimate.
+We show that this quantity, originally defined under a reference input distribution, can be estimated from samples drawn under auxiliary distributions by reweighting the model outputs.
+We derive the optimal sampling distribution that minimises the asymptotic variance of efficient estimators and demonstrate its impact on estimation.
+Beyond variance reduction, the framework also supports distributional sensitivity analysis through reverse importance sampling.
 1 Introduction
@@
 Since they are often costly to evaluate, handling uncertainty and performing sensitivity analysis is particularly critical.
-In the context of global sensitivity analysis (GSA), where model inputs are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the output [8, 4].
-To quantify this impact, a large variety of sensitivity indices has been proposed in the literature, each one associated with a different criterion.
-Some indices rely on higher-order moments [21, 24], others are based on divergence of measures between probability distributions [14, 18, 26], while others use contrast-based functionals or Cramér–von Mises-type distances [22, 33].
+In the context of global sensitivity analysis (GSA), where model inputs $\mathbf{X}=(X_{1},\dots,X_{k})$ are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the model output $Y$ [41, 42].
+To quantify this impact, a large variety of sensitivity indices have been proposed in the literature, each one associated with a different criterion.
+Some indices rely on higher-order moments [34, 36], others are based on discrepancy measures between probability distributions [4, 5, 10], while others use contrast-based functionals or Cramér–von Mises distance [20, 24].
 These indices allow for a refined view of the output distribution.
 Among all these tools, Sobol’ indices remain the most popular in practice.
-Originally introduced in [4], they are particularly well suited for scalar outputs, and can be generalized to vector-valued or functional outputs.
-Thanks to the Hoeffding decomposition [1], Sobol’ indices measure the contribution of each input to the output variance.
-As such, they are variance-based and only capture second-order effects.
+Originally introduced in [42], they are particularly well suited for scalar outputs, and can be generalised to vector-valued or functional outputs.
+Thanks to the Hoeffding decomposition [27], Sobol’ indices measure the contribution of each group of inputs to the output variance.
+The first-order Sobol’ indices quantify the main effects of each input, while the total-order Sobol’ indices also account for both main and interaction effects.
+Hence, they are variance-based measures that capture the contribution of inputs to the overall variability of the model output.
 In most practical applications, the model is a complex black-box code whose output distribution is unknown.
 Analytical formulas for Sobol’ indices are thus rarely available.
-As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis.
-Over the last decades, many estimation strategies have been proposed, which can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [30, 20]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [3], Random Balance Designs (RBD) [13], the Effective Algorithm for Sensitivity Indices (EASI) [17], and polynomial chaos expansions [15]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [29, 25]; and (iv) nonparametric estimators—including nearest-neighbor and kernel-based methods [11, 32, 36, 38, 35] and the rank-based estimator built on Chatterjee’s correlation coefficient [41, 38]—that require only a single input–output dataset.
+As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis, especially for the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$.
+Over the last decades, many estimation strategies have been proposed and can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [25, 35]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [9], Random Balance Designs (RBD) [47], the Effective Algorithm for Sensitivity Indices (EASI) [39], and polynomial chaos expansions [45]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [23, 28]; and (iv) nonparametric estimators including nearest-neighbor [17, 16, 7] and kernel-based methods [11, 21, 12], and the rank-based estimator [30, 22] built on Chatterjee’s correlation coefficient [8] which share the practical advantage of being applicable in the so-called given-data setting, where only a fixed input–output sample set is available and direct access to the black-box model is not possible.
 The theoretical properties of methods from categories (iii) and (iv) are now well established.
 In particular, for first-order Sobol’ indices, consistency and asymptotic normality have been proven for kernel estimators [12, 11], the Pick–Freeze estimator [28], nearest-neighbor methods [16], and rank-based estimators [22].
-Among them, the kernel estimator of [19] is efficient and achieves the asymptotic optimal variance bound for estimating the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$.However, it remains delicate to implement and computationally demanding.
-On the other hand, the rank-based estimator of [41] is much simpler to implement but sub-optimal and only suitable for first-order indices.
-A fundamental objective in Monte-Carlo estimation is to reduce the variance of the estimator without introducing bias.
-One of the most widely used techniques for variance reduction is importance sampling [2] (IS), which consists of modifying the sampling distribution to obtain a more efficient estimator of the quantity of interest.
-The variance of the IS estimator strongly depends on the choice of the sampling distribution.
+With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, as it requires two evaluations of the model for each realization of $X$, which is often prohibitive.
+When the context allows it, using this particular design can considerably simplify estimation, but the number of code calls often makes it prohibitive.
+By contrast, given-data methods are increasingly favored because they operate from a provided input–output sample only.
+This naturally raises the question of efficiency limits in the given-data setting.
+Among them, several estimators achieve the asymptotic efficiency lower bound.
+The expression of the efficiency bound for $\eta$ was first derived in [18] for a continuous input $X$, from the efficient influence function associated with the parameter.
+The same work also proposed an efficient estimator of a truncated version of $\eta$.
+Since then, several methods have been shown to be asymptotically efficient in the given-data setting.
+For first-order indices, the kernel estimator introduced in [11] achieves the asymptotic optimal variance bound for estimating $\eta$ (see Theorems 3.4 and 3.5).
+This result was initially established for first-order Sobol’ indices and later extended to higher-order ones under suitable regularity conditions [12] (see Theorems 4.1 and 4.2).
+These estimators, however, remain difficult to implement in practice.
+In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalized to attain the efficiency bound in [31] (Proposition 3.3).
+The question of efficiency bounds has been thoroughly discussed in [30].
+In this work, we mainly focus on these efficient estimators.
+Even though these estimators are already asymptotically optimal under the reference input distribution, their accuracy may degrade when the sample size is small or when the model response is highly nonlinear.
+In such cases, a large fraction of the simulation effort may be wasted on regions of the input space that contribute little to the Sobol’ quantity of interest, while the most informative regions are poorly explored.
+When the black-box model is still available, a natural way to improve efficiency is to modify the sampling scheme so that evaluations are more concentrated on informative regions.
+This motivates the use of importance sampling (IS) [29], one of the most classical and effective approaches: it modifies the sampling distribution to obtain a more accurate estimator of the target quantity.
+The variance achieved by IS strongly depends on how this sampling distribution is chosen.
 In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling.
-More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is often the main term requiring accurate estimation.
-We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution.
-We prove the existence of an optimal sampling distribution that can lead to zero asymptotic variance.
+More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is the most challenging term to estimate accurately.
+We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution, which makes it possible to characterize optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
 To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices.
-A notable exception is the contribution of [23], where the authors propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a quadrature scheme (Gauss–Hermite).
-Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte-Carlo simulation, thereby reducing numerical effort compared to classical procedures of Sobol’ indices estimation.
-However, their method does not address variance minimization explicitly.
-In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that allows estimation from alternative sampling distributions, and also offers computational benefits presented in Section 4.
-The paper is structured as follows, Section 2 introduces Sobol’ indices and the classical importance sampling framework.
+Two related contributions deserve to be mentioned.
+First, [2] propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a Gauss–Hermite quadrature scheme.
+Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte Carlo simulation, thereby reducing numerical effort compared to classical procedures.
+However, their method does not address variance minimisation explicitly.
+Second, [6] introduce a weighting method for estimating Sobol’ total effects under constrained input spaces, where an importance-type density ratio is used to correct Jansen’s estimator in non-Cartesian domains.
+In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that also extends naturally to distributional sensitivity analysis.
+Beyond variance reduction, the proposed importance sampling framework can also be exploited to assess the robustness of Sobol’ indices with respect to changes in the input distributions.
+This corresponds to the concept of reverse importance sampling [3, 1, 32], which exploits the fact that a single dataset can be reused to evaluate sensitivity measures under alternative probabilistic scenarios.
+Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [33].
+A post-processing importance sampling approach was also proposed in [46], where existing Monte Carlo samples are reweighted to estimate global sensitivity indices.
+Their method relies on a kernel-like weighting scheme to approximate conditional moments.
+The paper is structured as follows.
+Section 2 introduces Sobol’ indices and the classical importance sampling framework.
 Section 3 presents our main theoretical results.
-Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real test case.
+Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real-life test case.
 Section 5 concludes and outlines future research directions.
@@
 Definition of Sobol’ indices
-Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ a random vector independent from $\mathbf{X}$ with known distribution.
+Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square-integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ is a random vector independent of $\mathbf{X}$.
 In the context of sensitivity analysis, the so-called Sobol’ index $S_{u}$ is used to quantify the impact of a subset $u\subseteq\{1,\dots,k\}$ of the input variables on the output $Y$, and is defined as:
-$S_{u}=\frac{\mathrm{Var}\left(\mathbb{E}[Y\mid\mathbf{X}_{u}]\right)}{\mathrm{Var}(Y)}=\frac{\mathbb{E}\left[\mathbb{E}[Y\mid\mathbf{X}_{u}]^{2}\right]-\mathbb{E}[Y]^{2}}{\mathrm{Var}(Y)},$ (1)
-where:
-  • $\mathbf{X}_{u}=(X_{j})_{j\in u}$ denotes the subvector of $\mathbf{X}$ corresponding to the index set $u$,
-  • $\mathbf{X}_{\bar{u}}=(X_{j})_{j\in\bar{u}}$ is the complementary subvector, i.e., the remaining components of $\mathbf{X}$.
-As explained earlier, the main challenge for inference purposes lies in handling the conditional expectation $m_{u}(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}]$ in order to construct an estimator of
-$\eta_{u}=\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\left[Y\cdot m_{u}(\mathbf{X}_{u})\right]$ (2)
-In a given-data setting where an i.i.d.
-sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$ is available, optimal estimators of $\eta_{u}$ are known to be asymptotically Gaussian, with asymptotic variance achieving the efficiency bound [19] :
-$\sigma^{2}_{\mathrm{opt},u}(p)=4\mathbb{E}[m_{u}^{2}(X_{u})Y^{2}]-3\mathbb{E}\left[m_{u}^{4}(X_{u})\right]-\eta_{u}^{2}.$ (3)
-As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$, which drives the variance of Sobol’ index estimators.
+$S_{u}=\frac{\mathrm{Var}\!\left(m(\mathbf{X}_{u})\right)}{\mathrm{Var}(Y)},$ (1)
+where $m(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}=\mathbf{x}_{u}]$, with $\mathbf{X}_{u}=(X_{j})_{j\in u}$ the subvector of inputs corresponding to the index set $u$, and $\mathbf{X}_{-u}=(X_{j})_{j\notin u}$ its complementary subvector containing the remaining components of $\mathbf{X}$.
+One can develop the variance term as $\mathrm{Var}(m(\mathbf{X}_{u}))=\mathbb{E}[m^{2}(\mathbf{X}_{u})]-(\mathbb{E}[Y])^{2}$, and both expectations and variances can in practice be estimated through classical Monte Carlo estimators based on independent samples of $(Y,\mathbf{X})$.
+Hence, as recalled above, the main challenge for inference purposes lies in handling the conditional expectation $m(\mathbf{X}_{u})$ itself in order to construct an estimator of
+$\eta_{u}=\mathbb{E}\!\left[m^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\!\left[m(\mathbf{X}_{u})\,\mathbb{E}[Y\mid\mathbf{X}_{u}]\right]=\mathbb{E}[\,\mathbb{E}[Y\,m(\mathbf{X}_{u})\mid\mathbf{X}_{u}]\,]=\mathbb{E}\!\left[Y\,m(\mathbf{X}_{u})\right],$ (2)
+where the second equality follows from the law of total expectation.
+For an i.i.d.
+sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$, efficient estimators of $\eta_{u}$ are asymptotically Gaussian, with variance achieving the lower bound [18].
+Writing $\nu(\mathbf{X}_{u})=\mathrm{Var}(Y\mid\mathbf{X}_{u})$, the optimal asymptotic variance is given by:
+$\sigma^{2}_{\mathrm{opt},u}=4\,\mathbb{E}\!\left[m^{2}(\mathbf{X}_{u})\,\nu(\mathbf{X}_{u})\right]+\mathrm{Var}\!\left(m^{2}(\mathbf{X}_{u})\right).$ (3)
+As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$.
 For both theoretical and practical reasons, it is often beneficial to modify the original distribution of the input variables.
-In the importance sampling framework, briefly introduced in the next subsection, we consider the following setting: the input vector $\mathbf{X}$ is initially distributed according to a reference distribution $p$, under which the components are assumed to be dependent, although independence can also be assumed for analytical convenience.
-We then consider the use of an alternative sampling distribution $q$, potentially introducing a dependency structure.Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$ and using samples $\mathbf{X}\sim q$ instead of $\mathbf{X}\sim p$.
-To avoid any ambiguity, we explicitly indicate the distribution of the input $\mathbf{X}$ in the expectation notation.
-For example, $\mathbb{E}_{p}[\cdot]$ denotes the expectation when $\mathbf{X}$ is distributed according to the density $p$.
-The distribution $W$ is fixed and remains unchanged throughout.
-The distribution of the output $Y$ is only affected through changes in the distribution of the input.
-Importance sampling (IS)
-A classical variance-reduction technique.
-To estimate an expectation of the form $\mathbb{E}_{p}[f(\mathbf{X})]$, one rewrites it under a distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$ where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ is the likelihood ratio.
-To ensure the estimator is unbiased, the support of $q$ must include the support of $f(\mathbf{x})\,p(\mathbf{x})$.
-Given an i.i.d sample $\mathbf{X}^{(1)},...,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
+Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$.
+In the importance sampling framework, briefly recalled below, the input vector $\mathbf{X}$ is initially drawn from a reference distribution $p$, under which the components may be dependent, although independence can be assumed for analytical convenience.
+We then introduce an alternative sampling distribution $q$, possibly with a different dependency structure.
+To avoid ambiguity, expectations are indexed by the underlying distribution: for instance, $\mathbb{E}_{p}[\cdot]$ denotes expectation under $p$.
+The distribution of $\mathbf{W}$ is fixed and remains unchanged, and the output $Y$ is affected only through changes in the distribution of $\mathbf{X}$.
+Importance sampling
+This classical variance-reduction method estimates expectations of the form $\mathbb{E}_{p}[f(\mathbf{X})]$ by rewriting them under another distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$, where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ denotes the likelihood ratio.
+To ensure unbiasedness, the support of $q$ must include that of $f(\mathbf{x})\,p(\mathbf{x})$.
+Given an i.i.d.
+sample $\mathbf{X}^{(1)},\dots,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
 $\hat{I}^{\mathrm{IS}}=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{X}^{(i)})\,w(\mathbf{X}^{(i)}).$ (4)
-This estimator achieves zero variance if and only if $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost-everywhere.
-This ideal distribution is not usable in practice because it depends on the unknown quantity to be estimated.
-However, adaptive procedures exist to approximate it, such as non-parametric methods [7] or parametric approaches like the cross-entropy method [34, 12].
-It is worth noting that the effectiveness of IS [9] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large.
+This estimator achieves zero variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere.
+This ideal distribution cannot be used in practice because it depends on the unknown quantity to be estimated.
+However, adaptive procedures exist to approximate it, such as nonparametric methods [49] or parametric approaches like the cross-entropy method [38, 14].
+The efficiency of IS [37] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large.
 Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values.
-The next section presents our main contributions by extending the classical IS framework to characterize the optimal sampling distribution $q$ for estimating $\eta_{u}$ and improving Sobol’ indices estimation.
+The next section extends the classical IS framework to characterize the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
 3 Main results
 3.1 Input distribution change via importance sampling
-We consider a modification of the input distribution, applied exclusively to the random vector $\mathbf{X}\in\mathbb{R}^{k}$.
-Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$, for any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\textbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x})$ its marginal density and introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
-$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}),$ (5)
-where $q_{u}(\mathbf{x}_{u})$ and $q_{\bar{u}}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional densities, respectively. The corresponding importance weights are then given by:
-$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\cdot p_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\cdot w_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}).$ (6)
-Lemma
-Let $q=q_{u}\cdot q_{\bar{u}}$ be any importance sampling distribution on the input $\mathbf{X}$, define
-$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot Y.$ (7)
+We introduce the importance sampling framework by allowing a modification of the input distribution of the random vector $\mathbf{X}\in\mathbb{R}^{k}$, while keeping the noise variable $\mathbf{W}$ fixed.
+Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$.
+For any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\mathbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x}_{u})$ its marginal density.
+We introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
+$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\,q_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$ (5)
+where $q_{u}(\mathbf{x}_{u})$ and $q_{-u\mid u}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional components of $q$. The corresponding importance weights are then given by
+$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\,p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\,q_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\,w_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}).$ (6)
+Lemma 3.1
+Let $q=q_{u}\,q_{-u\mid u}$ be any importance sampling distribution on the input $\mathbf{X}$, and define
+$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,Y.$ (7)
 Then:
-$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}).$ (8)
+$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\,m(\mathbf{X}_{u}).$ (8)
 In particular, the following identity holds:
-$\eta_{u}=\mathbb{E}_{p}\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{2}\right]=\mathbb{E}_{q}\left[Z_{u}\cdot\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
-Proof
-See Appendix
-This identity provides a reformulation of $\eta_{u}$ under the sampling distribution $q$ and can be consistently estimated using the auxiliary variable $Z_{u}$.
-We will show next that a suitable choice of $q$ leads to a reduced variance of the corresponding estimator.
+$\eta_{u}=\mathbb{E}_{p}\!\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\!\left[(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\!\left[Z_{u}\,\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
+The proof is given in the Appendix.
+In the following, this identity is used to express $\eta_{u}$ under $q$ and to show that an appropriate choice of $q$ can reduce the variance of its estimator.
 3.2 Optimal asymptotic variance expression using importance sampling
-We now investigate the case where the reference distribution of the input vector $\mathbf{X}$ can be freely modified, while the distribution of the noise variable $W$ remains fixed.
-We define $\phi(\mathbf{X})=\sqrt{\mathbb{E}[Y^{2}|\mathbf{X}]}$, noting that $\phi$ is model-specific and independent of the sampling distribution; its dependence on $p$ or $q$ does not need to be made explicit in the notation.
+We define $\phi(\mathbf{x})=\sqrt{\mathbb{E}[Y^{2}\mid\mathbf{X}=\mathbf{x}]}$, noting that $\phi$ is model-specific and does not depend on the choice of sampling distribution.
 Under $p$, the optimal asymptotic variance for estimating $\eta_{u}$ is given by
-$\sigma^{2}_{\mathrm{opt},u}(p)=4\,\mathbb{E}_{p}\left[m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]-3\,\mathbb{E}_{p}\left[m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$ (10)
-We now derive the expression of the asymptotic variance under an arbitrary auxiliary distribution $q$, and compare it to the reference case.
-proposition
-Let $q(\mathbf{x})$ be a sampling density over $\mathbf{X}$, and let $Z_{u}$ be defined as in Lemma 1. Then the optimal asymptotic variance is given by
-$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$ (11)
-Using Lemma 1 and substituting the expressions of $Z_{u}$, this variance can be rewritten under $p$ as
-$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]$ (12)
-$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
-This expression highlights how the asymptotic variance depends explicitly on the choice of the sampling distribution $q$ through the weights $w_{u}$ and $w_{\bar{u}}$, and through the structure of $Z_{u}$.
-Optimizing $q$ can therefore lead to substantial variance reduction.
+$\sigma^{2}_{\mathrm{opt},u}(p)=\mathbb{E}_{p}\Big[4\,m^{2}(\mathbf{X}_{u})\,\nu_{p}(\mathbf{X}_{u})+m^{4}(\mathbf{X}_{u})\Big]-\eta_{u}^{2},$ (10)
+where the conditional variance $\nu_{p}$ is given by
+$\nu_{p}(\mathbf{X}_{u}):=\mathrm{Var}_{p}(Y\mid\mathbf{X}_{u})=\mathbb{E}_{p}[Y^{2}\mid\mathbf{X}_{u}]-m^{2}(\mathbf{X}_{u})=\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-m^{2}(\mathbf{X}_{u}),$
+We next extend this result to an auxiliary distribution $q$ and compare it with the reference case.
+Proposition 3.2
+Let $Z_{u}$ be defined as in Lemma 3.1. Then, for the sampling distribution $q$, the optimal asymptotic variance is given by
+$\boxed{\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.}$ (11)
+Substituting the expressions of $Z_{u}$, $\sigma^{2}_{\mathrm{opt},u}(q)$ can be equivalently expressed under $p$ as
+$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\Big[w_{u}(\mathbf{X}_{u})\big(4\,m^{2}(\mathbf{X}_{u})\,\nu_{q}(\mathbf{X}_{u})+m^{4}(\mathbf{X}_{u})\big)\Big]-\eta_{u}^{2},$ (12)
+where
+$\nu_{q}(\mathbf{X}_{u})=\mathbb{E}_{p}\big[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\big]-m^{2}(\mathbf{X}_{u}).$
+This expression shows that the asymptotic variance depends explicitly on $q$ through the weights $w_{u}$ and $w_{-u\mid u}$, and can be reduced by optimising them, as will be shown in the next section.
 3.3 Optimal distribution: sequential construction
-We aim to minimize the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimizing the components of the sampling distribution $q$ in a sequential manner.
-Our strategy proceeds in two steps.
-First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$ that minimizes the variance contribution given $\mathbf{x}_{u}$.
-This optimal conditional density is then substituted into the variance expression.
-In the second step, the resulting objective is minimized with respect to the marginal $q_{u}$.
-To characterize the optimal distribution, we rely on a key variational lemma recalled below.
-Lemma
-Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$.
-Then, for any $g\in\mathcal{D}_{k}$, the value $q\overset{ae}{=}g$ uniquely minimizes the functional
-$q\mapsto\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x}\quad,\quad q\in\mathcal{D}_{k}.$
-Proof
-Let $q\in\mathcal{D}_{k}$. By Jensen’s inequality applied to the convex function $t\mapsto t^{2}$, we obtain:
-[…]
-Where equality is achieved if and only if $q=g$ almost everywhere.
-First step: optimization of $q_{\bar{u}\mid u}$
-We begin by optimizing the conditional component $q_{\bar{u}\mid u}$ of $q$, assuming that $q_{u}$ is fixed.
-From Proposition 3.2, the conditional distribution influences the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ in eq(3.6) only through the term
-$\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
-Minimizing this expectation is therefore sufficient, since it is the only component of the variance that depends on $q_{\bar{u}\mid u}$.
-Applying Lemma 2 conditionally on $\mathbf{X}_{u}$ gives the optimal conditional density:
-$q^{*}_{\bar{u}\mid u}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})=\frac{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi(\mathbf{X})}{\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}\right]},$ (13)
-where the denominator ensures normalization.
-With this choice, the variance contribution when $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$ becomes $\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{\bar{u}\mid u}\sim p_{\bar{u}}$ (i.e., $w_{\bar{u}\mid u}=1$).
+We aim to minimise $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimising the components of $q$ in a sequential manner.
+The procedure involves two steps.
+First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})$ that minimises the variance contribution given $\mathbf{x}_{u}$, which is then substituted into the variance expression.
+In the second step, the resulting objective is minimised with respect to $q_{u}$.
+To characterize the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations.
+Following the classical Lagrangian minimisation framework, one can formally derive the stationary condition satisfied by $q^{*}$ under the normalisation constraint.
+In general, verifying that this stationary point corresponds to a minimum would require checking second-order optimality conditions.
+However, as mentioned by Owen [34, End Notes, p.
+38], the Cauchy–Schwarz inequality is sufficient to establish global minimality in this setting.
+For simplicity, we rely on the following key variational lemma, a standard result in importance sampling theory.
+Lemma 3.3
+Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$. Then, for any $g\in\mathcal{D}_{k}$, the functional
+$J(q)=\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x},\qquad q\in\mathcal{D}_{k}.$
+is uniquely minimised by $g$ a.e.
+Proof 1
+We apply Cauchy–Schwarz inequality:
+$J(g)=1=\left(\int g(\mathbf{x})\,d\mathbf{x}\right)^{2}=\left(\int\frac{g(\mathbf{x})}{q(\mathbf{x})}\,q(\mathbf{x})\,d\mathbf{x}\right)^{2}\leq\int\left(\frac{g(\mathbf{x})}{q(\mathbf{x})}\right)^{2}q(\mathbf{x})\,d\mathbf{x}=J(q),$
+Equality in the Cauchy–Schwarz inequality holds if and only if the argument of the convex function is constant almost everywhere with respect to $q(\mathbf{x})\,d\mathbf{x}$.
+That is, there exists a constant $c>0$ such that $g(\mathbf{x})/q(\mathbf{x})=c$ a.e.
+Since both $g$ and $q$ integrate to one, necessarily $c=1$, and hence $q=g$ almost everywhere.
+First step: optimisation of $q_{-u\mid u}$
+We optimise the conditional component $q_{-u\mid u}$ of $q$ for a fixed $q_{u}$.
+Since $\sigma^{2}_{\mathrm{opt},u}(q)$ in Eq.
+(12) depends on $q_{-u\mid u}$ only through the term $\mathbb{E}_{p}\!\left[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right],$ which appears in the definition of $\nu_{q}(\mathbf{X}_{u})$, minimising this conditional expectation is sufficient.
+Applying Lemma 3.3 conditionally on $\mathbf{X}_{u}$ yields the optimal conditional density
+$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,\phi(\mathbf{x})}{\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]},$ (13)
+where the denominator ensures normalisation.
+With this choice, the variance contribution when $q_{-u\mid u}=q^{*}_{-u\mid u}$ becomes $\big(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\big)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{-u}\mid\mathbf{X}_{u}\sim p_{-u\mid u}$ (i.e., $w_{-u\mid u}=1$).
 A full derivation is provided in the Appendix.
-Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality confirms that
-$\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
-which guarantees that the optimal conditional distribution $q^{*}_{\bar{u}\mid u}$ systematically reduces this part of the variance.
+Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality implies
+$\big(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\big)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
+showing that $q^{*}_{-u\mid u}$ indeed reduces this part of the variance.
 Second step: optimisation of the marginal $q_{u}$
-We now fix the component $q_{\bar{u}\mid u}$, and seek to minimize the asymptotic variance over the marginal $q_{u}$.
-For any such fixed conditional, the sampling density is of the form $q=q_{u}\cdot q_{\bar{u}\mid u}$, and the corresponding variance reads
-$\sigma_{\operatorname{opt},u}^{2}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
-where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{\bar{u}\mid u}$. The optimal marginal minimizing this expression is given by
-$q_{u}^{*}(\mathbf{x}_{u})=\frac{p(\mathbf{x}_{u})\,\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\operatorname{opt},u}(q)=\left(\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
+We now fix $q_{-u\mid u}$, and seek to minimise the asymptotic variance over $q_{u}$.
+For any such fixed conditional, the sampling density is of the form $q=q_{u}\,q_{-u\mid u}$, and the corresponding variance reads
+$\sigma_{\mathrm{opt},u}^{2}(q)=\mathbb{E}_{p}\!\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
+where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{-u\mid u}$. The optimal marginal minimising this expression is given by
+$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,\sqrt{S_{q_{-u\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\!\left[\sqrt{S_{q_{-u\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q_{u}^{*})=\left(\mathbb{E}_{p}\!\left[\sqrt{S_{q_{-u\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
 We distinguish below two particular cases:
-  • Case A – Reference conditional: $q_{\bar{u}\mid u}=p_{\bar{u}\mid u}$
-    $S_{A}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-3\,m_{u}^{4}(\mathbf{x}_{u}).$
-  • Case B – Optimal conditional: $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$
-    $S_{B}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}-3\,m_{u}^{4}(\mathbf{x}_{u}).$
-This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, shown in step one. The full chain of inequalities then reads
-$\sigma^{2}_{\operatorname{opt},u}(q^{*})\leq\sigma^{2}_{\operatorname{opt},u}(q_{u}^{*}\cdot p_{\bar{u}\mid u})\leq\sigma^{2}_{\operatorname{opt},u}(p),$
-this follows from Jensen’s inequality:
-$\left(\mathbb{E}_{p}[\sqrt{S_{B}(\mathbf{X}_{u})}]\right)^{2}\leq\left(\mathbb{E}_{p}[\sqrt{S_{A}(\mathbf{X}_{u})}]\right)^{2}\leq\mathbb{E}_{p}[S_{A}(\mathbf{X}_{u})].$
-That is, joint optimisation in Case B achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
-3.4 Zero-variance optimal distribution
-In the idealised case where the function $f$ is fully deterministic with all inputs controllable (i.e., there is no external noise variable $W$ and $Y=f(\mathbf{X})$), it becomes theoretically possible to achieve zero-variance in the estimation of $\eta_{u}$.
-This setting can be interpreted as a special case of the general framework discussed earlier.
-We still proceed in two steps, but expressions simplify significantly.
+  • Case A – Reference conditional: $q_{-u\mid u}=p_{-u\mid u}$
+    $S_{A}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{p}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u})$
+  • Case B – Optimal conditional: $q_{-u\mid u}=q^{*}_{-u\mid u}$
+    $S_{B}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{q^{*}}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\big(\mathbb{E}_{p}\!\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}\right]\big)^{2}-3m^{4}(\mathbf{x}_{u}),$
+Both $S_{A}(\mathbf{x}_{u})$ and $S_{B}(\mathbf{x}_{u})$ are non-negative since they are built from conditional variances $\nu_{p}(\mathbf{X}_{u})$ and $\nu_{q}(\mathbf{X}_{u})$.
+This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, as shown in the first step. The full chain of inequalities then reads
+$\sigma^{2}_{\mathrm{opt},u}(q^{*})\leq\sigma^{2}_{\mathrm{opt},u}(q_{u}^{*}\,p_{-u\mid u})\leq\sigma^{2}_{\mathrm{opt},u}(p),$
+which follows from Jensen’s inequality:
+$\left(\mathbb{E}_{p}\!\left[\sqrt{S_{B}(\mathbf{X}_{u})}\right]\right)^{2}\leq\left(\mathbb{E}_{p}\!\left[\sqrt{S_{A}(\mathbf{X}_{u})}\right]\right)^{2}\leq\mathbb{E}_{p}\!\left[S_{A}(\mathbf{X}_{u})\right].$
+Joint optimisation in Case B therefore achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
+3.4 Zero variance optimal distribution
+In the idealised deterministic case, where all inputs are controllable and no external noise variable $\mathbf{W}$ is present ($Y=f(\mathbf{X})$), it is theoretically possible to achieve zero variance in estimating $\eta_{u}$.
+This case is a direct simplification of the general framework, with the same two-step procedure but simpler expressions.
+We still proceed in two steps, but the expressions simplify significantly.
 In particular, the asymptotic variance now takes the form
-$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y^{2}\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
-Applying Lemma 2 sequentially yields:
-$q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (16)
+$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\!\left[w_{u}(\mathbf{X}_{u})\left(4\,m^{2}(\mathbf{X}_{u})\,\mathbb{E}_{p}\!\left[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,Y^{2}\,\big|\,\mathbf{X}_{u}\right]-3\,m^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
+Applying Lemma 3.3 sequentially yields
+$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,|f(\mathbf{x})|}{\mathbb{E}_{p}\!\left[|f(\mathbf{X})|\,\big|\,\mathbf{X}_{u}=\mathbf{x}_{u}\right]},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,\sqrt{S_{C}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\!\left[\sqrt{S_{C}(\mathbf{X}_{u})}\right]},$ (16)
+where
+$S_{C}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\Big(\mathbb{E}_{p}\!\left[|f(\mathbf{X})|\,\big|\,\mathbf{X}_{u}=\mathbf{x}_{u}\right]\Big)^{2}-3\,m^{4}(\mathbf{x}_{u}).\\$
+Hence, the optimal joint sampling distribution is $q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q_{-u\mid u}^{*}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$ from which we obtain
+$\sigma^{2}_{\mathrm{opt},u}(q^{*})\leq\sigma^{2}_{\mathrm{opt},u}(p).$
+In general, this distribution does not yield a zero variance estimator, since a direct plug-in calculation shows
+$\sigma^{2}_{\mathrm{opt},u}(q^{*})=\left(\mathbb{E}_{p}\!\left[\sqrt{S_{C}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}>0.$
+Remark.
+When the model $f(\mathbf{x})$ may take both positive and negative values, the theoretical zero variance density derived above is no longer a valid probability distribution.
+A practical alternative consists in separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution.
+This idea, sometimes referred to as positivisation in the importance sampling literature [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
+We decompose $f$ into its positive and negative parts:
+$f_{+}(\mathbf{x})=\max(f(\mathbf{x}),0),\qquad f_{-}(\mathbf{x})=\max(-f(\mathbf{x}),0),\qquad f(\mathbf{x})=f_{+}(\mathbf{x})-f_{-}(\mathbf{x}).$
+Two auxiliary densities $q_{+}$ and $q_{-}$ are then defined to be positive wherever $pf_{+}$ and $pf_{-}$ are positive, respectively.
+Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, preserving unbiasedness while reducing variance in practice.
+If $f(\mathbf{x})\geq 0$ for all $\mathbf{x}$, then $\mathbb{E}_{p}[|f(\mathbf{X})|\mid\mathbf{X}_{u}]=m(\mathbf{X}_{u})$ and $S_{C}(\mathbf{x}_{u})=m^{4}(\mathbf{x}_{u})$, so that
+$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{m(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (17)
 The resulting fully adapted sampling density is
-$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\cdot q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m_{u}(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (17)
-Hence, the density $q^{*}(\mathbf{x})$ achieves a zero-variance estimator in theory.
-However, this ideal result is not achievable in practice, as the optimal distribution depends on the unknown quantity $\eta_{u}$ that we aim to estimate.
+$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
+In theory, the density $q^{*}(\mathbf{x})$ yields a zero variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it.
+Adaptive procedures can be employed, including nonparametric methods [49] and parametric ones such as the cross-entropy method [38, 14].
 4 Numerical Experiments
-4.1 Analytical Illustration of Variance Reduction via IS
-In this section, we illustrate the effectiveness of an IS transformation in reducing the variance $\eta_{u}$ estimators. As a benchmark example, we consider the classical Sobol’ g-function [8]:
+4.1 Analytical illustration of variance reduction
+In this section, we illustrate the effectiveness of an importance sampling transformation in reducing the variance of $\eta_{u}$ estimators.
+As a benchmark example, we consider the classical Sobol’ g-function [41]:
 $g(X_{1},\dots,X_{k})=\prod_{i=1}^{k}\frac{|4X_{i}-2|+a_{i}}{1+a_{i}},$ (19)
-where $X_{i}\sim\mathcal{U}[0,1]$ are assumed to be independent inputs.
-This function is widely used in sensitivity analysis due to its non-linear, non-monotonic nature and the availability of analytical Sobol’ indices.
-We consider a 3-dimensional case $(k=3)$ with $a_{i}=i$, and focus on estimating the second-order Sobol’ index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$.
+where $X_{i}\sim\mathcal{U}[0,1]$ are independent inputs.
+This function is a standard benchmark in sensitivity analysis due to its nonlinear, non-monotonic structure and the availability of analytical Sobol’ indices.
+We consider the 3-dimensional case $(k=3)$ with $a_{i}=i$ and focus on the second-order index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$.
+For this part, one could in principle rely on kernel estimators [18, 11, 12], but since all quantities are analytically tractable, we rely on their closed forms and focus on the impact of optimal importance densities on variance reduction.
 The third input $X_{3}$ is treated as an uncontrolled variable $W$.
-We consider the case A in (3.3), where $\bar{u}=\emptyset$.
-In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically, including the optimal sampling density.
-Specifically, the optimal density $q^{\star}$ to estimate $\eta_{12}$ is given by
+We consider Case A in Section 3.3, where $u=\emptyset$.
+In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically.
+Specifically, the optimal density $q^{*}$ to estimate $\eta_{\{1,2\}}$ is given by
 $q^{*}(x_{1},x_{2})=\frac{p(x_{1},x_{2})\cdot\sqrt{S(x_{1},x_{2})}}{\mathbb{E}_{p}[\sqrt{S(X_{1},X_{2})}]},\quad\text{with}\quad S(x_{1},x_{2})=4\,m^{2}(x_{1},x_{2})\,\phi^{2}(x_{1},x_{2})-3\,m^{4}(x_{1},x_{2}).$
@@
 $S(x_{1},x_{2})=\frac{9}{8}\cdot m^{4}(x_{1},x_{2}),\quad\text{so that}\quad q^{*}(x_{1},x_{2})=\frac{m^{2}(x_{1},x_{2})}{\eta_{\{1,2\}}}.$
-We first plot the shape of $q^{*}$ over $[0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
+Remark.
+Although our example involves an uncontrolled variable $W=X_{3}$, its particular structure satisfies $\phi^{2}(x_{u})\propto m^{2}(x_{u})$, so that the function $S_{A}(x_{u})$ simplifies to $S_{A}(x_{u})=c\,m^{4}(x_{u})$ for some constant $c>0$.
+As a result, $q_{u}^{*}(x_{u})$ retains the same analytical form as in the noise-free case, namely $q_{u}^{*}(x_{u})\propto m^{2}(x_{u})$.
+This coincidence is specific to the g-function and does not hold in general models.
+We first plot the shape of $q^{*}$ over $0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
 $q_{t}(x_{1},x_{2})=(1-t)\,p(x_{1},x_{2})+t\,q^{*}(x_{1},x_{2}),\quad t\in[0,1].$
@@
 $\mathrm{CV}_{\mathrm{opt}}(t)=\frac{\sigma_{\mathrm{opt},u}(q_{t})}{\eta_{\{1,2\}}},\quad\text{with}\quad\sigma^{2}_{\mathrm{opt},u}(q_{t})=\mathbb{E}_{p}\left[\frac{S(X_{1},X_{2})}{q_{t}(X_{1},X_{2})}\right]-\eta_{\{1,2\}}^{2}.$
-Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte-Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
-(a) Optimal density $q^{\star}(x_{1},x_{2})$
+Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
+(a) Optimal density $q^{*}(x_{1},x_{2})$
 (b) Coefficient of variation as a function of $t$
-Figure 1: Left: optimal sampling density $q^{\star}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{\star}$.
-Parametric optimization
+Figure 1: Left: optimal sampling density $q^{*}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{*}$.
+Parametric optimisation
 An alternative strategy consists of introducing a parametric family of auxiliary sampling densities for $(X_{1},X_{2})$, defined by
@@
 where the shape parameters $\alpha_{i},\beta_{i}$ are positive real numbers, i.e., $\alpha_{i},\beta_{i}\in\mathbb{R}_{+}^{*}$ for $i=1,2$.
-We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform distribution at $(\alpha,\beta)=(1,1)$.
-The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimizes the asymptotic variance, namely
+We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform one at $(\alpha,\beta)=(1,1)$.
+The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimises the asymptotic variance, namely
 $(\alpha^{*},\beta^{*})=\arg\min_{(\alpha,\beta)\in(\mathbb{R}_{+}^{*})^{2}}\sigma^{2}_{\mathrm{opt},u}(q_{\alpha,\beta}).$
@@
 Figure 2: Asymptotic variance $\sigma_{\mathrm{opt}}^{2}(\alpha,\beta)$ over the Beta parameter grid (symmetric case) displayed on a logarithmic scale.
+The optimal configuration is found numerically around $(\alpha,\beta)\approx(0.7,\,0.7)$.
+This yields a minimal variance of approximately $0.3368$, corresponding to a reduction of about $60\%$ compared to the uniform case $(\alpha=\beta=1)$.
+The optimal parameters are marked in red on the heatmap.
+4.2 Sampling and estimation example
+To move closer to a realistic setting, we now consider an estimation scenario where the target quantity $\eta_{\{1,2\}}$ is unknown and must be estimated from the observed data.
+We work with the same Sobol’ g-function and replace analytical quantities by Monte Carlo estimates while using the oracle $m(\mathbf{x})$ to isolate the impact of $q^{*}$.
+We compare two strategies:
+  • Reference case: samples drawn from the uniform input $p$. In this case, we estimate
 […]
+  • Importance sampling case: samples are drawn from an empirical estimate of the optimal density $q^{*}\propto m^{2}(x_{1},x_{2})\,\mathbf{1}_{[0,1]^{2}}(x)$, as shown in the first part, denoted by $\widehat{q}_{\theta}$.
+The corresponding estimator reads
+[…]
+The Beta parametric family identified numerically in the first example can still be used as a simple baseline.
+In the following, however, we focus on a more flexible parametric approximation of $q^{*}$.
+Since $q^{*}(x_{1},x_{2})\propto m^{2}(x_{1},x_{2})$, its evaluation requires calls to the model $g(x_{1},x_{2},X_{3})$ to compute the conditional mean $m(x_{1},x_{2})$; consequently, the generation of samples approximately following $q^{*}$ depends on model evaluations.
+A preliminary sample $(\mathbf{x}^{(i)})_{i=1}^{N_{\mathrm{fit}}}$ is obtained by a simple acceptance–rejection algorithm [19]:
+$\mathbf{x}^{(i)}\sim\mathrm{Unif}([0,1]^{2}),\qquad u^{(i)}\sim\mathrm{Unif}(0,M),\qquad\text{accept if }u^{(i)}\leq m^{2}(\mathbf{x}^{(i)}),$
+with $M=\sup_{[0,1]^{2}}m^{2}(\mathbf{x})$. The accepted points follow $q^{*}$ up to normalisation and are used to fit $q_{\theta}$.
+Figure 3: Samples generated via rejection sampling following $q^{*}$.
+To preserve the support $0,1]^{2}$ while retaining flexibility, we use a logit–Gaussian mixture model (logit–GMM), where $\mathrm{logit}(x)=\log\!\tfrac{x}{1-x}$ and $\mathrm{logistic}(z)=(1+e^{-z})^{-1}$.
+Each pilot point $\mathbf{x}=(x_{1},x_{2})$ is mapped to the unbounded space $\mathbf{z}=(z_{1},z_{2})$ through $z_{j}=\mathrm{logit}(x_{j})$, and a Gaussian mixture is then fitted:
+$q_{Z}(\mathbf{z};\theta)=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(\mathbf{z};\mu_{k},\Sigma_{k}),\qquad\pi_{k}>0,\quad\sum_{k=1}^{K}\pi_{k}=1,$
+with parameters $\theta=\{\pi_{k},\mu_{k},\Sigma_{k}\}_{k=1}^{K}$.
+By the change-of-variables formula, this induces a properly normalised expression on $0,1]^{2}$:
+$q_{\theta}(\mathbf{x})=q_{Z}\!\bigl(\mathrm{logit}(\mathbf{x});\theta\bigr)\prod_{j=1}^{2}\frac{1}{x_{j}(1-x_{j})},\qquad\mathbf{x}\in(0,1)^{2}.$
+[…]
+[…]
+This experiment mimics a practical learning-based IS setting, where the optimal distribution $q^{*}$ is unknown and approximated from data, leading to noticeable variance reduction.
 4.3 Distributional Sensitivity via Reverse Importance Sampling
-We study the sensitivity of the quantity
-$\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m_{u}^{2}(\mathbf{X}_{u})\right],$
-with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$.
-This falls into the framework of reverse importance sampling; this method, originally referred to as importance sampling sensitivity analysis [6], is now often described as a reverse importance sampling mechanism [27] and exploits the fact that a single set of observations can be reused to estimate responses under multiple reference distributions.
-Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [5].
+We study the sensitivity of the quantity $\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m^{2}(\mathbf{X}_{u})\right],$ with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$.
+This falls into the framework of reverse importance sampling.
 The rank estimator [22], based on the sequential correlation of outputs, takes as input a sequence $Y^{(i)}$ sorted with respect to the associated inputs $\mathbf{X}_{u}^{(i)}$, and converges to $\eta_{u}$ via the average of consecutive products.
-While initially defined under the distribution $p$, this estimator can be naturally extended to estimate $\eta_{u}(\theta)$ under any distribution $p_{\theta}$, using the results from Section 3.1, the convergence to $\eta_{u}(\theta)$ is guaranteed.
-For each $\theta$, we define the reweighted output
-$Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\cdot\frac{p_{\theta,\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}\cdot Y,$ (19)
-so that the rank-based estimator becomes
-$\hat{\eta}^{\text{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)}.$ (20)
+By using the results from Section 3.1, the same estimator can be extended to compute $\eta_{u}(\theta)$ under any distribution $p_{\theta}$.
+This formulation allows extending the rank-based estimator of [22] to any perturbed distribution $p_{\theta}$ through reverse importance sampling and remains consistent.
+Its asymptotic behavior follows directly from the consistency result established in Proposition 3.2 of [22].
+We note that in this section we work with the rank estimator (and not the lagged–rank estimator), which is not asymptotically efficient in terms of variance.
+Since our purpose here is only to study the sensitivity of the Sobol’ indices under different input distributions, the optimal densities derived in Section 3 are not involved.
+For completeness, we nevertheless derive in Appendix B the optimal sampling densities associated with the rank estimator itself.
+For each $\theta$, we define the reweighted output and the rank-based estimator becomes
+$\widehat{\eta}^{\mathrm{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)},\qquad Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\,\frac{p_{\theta,-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})}{p_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})}\,Y.$ (20)
 Hence, from a single dataset generated under $p$, one can efficiently explore the sensitivity of $\eta_{u}$ (and thus the associated Sobol’ index $S_{u}$) across the entire family of distributions $p_{\theta}$, without performing any additional evaluations of the black-box model.
@@
 Physical model
-[…]
+We consider a black-box simulator based on the IGLOO2D code developed at ONERA [48], which returns the ice-accretion thickness along an airfoil profile.
+Each run of the simulator takes as input a vector of four physical variables $\mathbf{X}=(X_{1},X_{2},X_{3},X_{4})\in\mathbb{R}^{4}$, initially sampled independently from uniform distributions defined over the intervals $\mathbf{x}_{\mathrm{low}}=(3.5,\;65,\;261.15,\;4\times 10^{-4})$ and $\mathbf{x}_{\mathrm{up}}=(4.5,\;75,\;265.15,\;6\times 10^{-4})$, which respectively correspond to $X_{1}$ (angle of attack, AoA), $X_{2}$ (freestream velocity, $U_{\infty}$), $X_{3}$ (ambient temperature, $T_{\mathrm{inf}}$), and $X_{4}$ (liquid water content, LWC).
+A design of experiments of size $N=2500$ has been generated, yielding input samples $\mathbf{X}^{(i)}\in\mathbb{R}^{4}$ for $i=1,\dots,N$, and corresponding outputs $Y^{(i)}\in\mathbb{R}$, which represent the accretion thickness at a given location along the airfoil.
+All input variables $X_{j}$ are first normalised to the unit interval via $X_{j}^{\mathrm{std}}:=(X_{j}-x_{\mathrm{low},j})/(x_{\mathrm{up},j}-x_{\mathrm{low},j})$.
+To explore the sensitivity of this output to changes in the input distributions, we consider alternative input distributions $p_{\theta_{j}}(X_{j})$ built from rescaled $\operatorname{Beta}(\theta_{j})$ marginals on the same support, where varying $\theta_{j}=(\alpha_{j},\beta_{j})$ allows smooth deviations from the uniform case $(1,1)$.
+For any configuration $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$, the corresponding Sobol’ index $S_{u}(\boldsymbol{\theta})$ or contrast $\eta_{u}(\boldsymbol{\theta})$ can be estimated from the same dataset by importance weighting, using the factor
 $w_{\boldsymbol{\theta}}(\mathbf{X})=\prod_{j=1}^{4}\frac{\operatorname{Beta}_{\theta_{j}}(X_{j}^{\mathrm{std}})}{\operatorname{Unif}[0,1](X_{j}^{\mathrm{std}})}.$ (21)
-With the rank estimator in eq(4.3), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$.
+With the rank estimator in Eq.
+(20), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$.
 Two perturbation strategies are considered: (i) marginal perturbation, where only the pair $\theta_{j}=(\alpha_{j},\beta_{j})$ associated with a single input $X_{j}$ is varied while the remaining marginals remain uniform; and (ii) global perturbation, where several or all input marginals are varied simultaneously.
@@
 All curves are computed from the same baseline sample ($N=2500$).
-These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, with markedly different behaviors across the four inputs.
-Overall, these confirm that the proposed given-data IS framework offers a powerful and cost-effective mean to perform detailed distributional local sensitivity analyses of Sobol’ indices.
+These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, which varies through the two parameters $\alpha_{j}$ and $\beta_{j}$ of the Beta distribution, with different behaviours across the four inputs.
+The strongest variations are observed for the pair $(\alpha_{4},\beta_{4})$, which is expected since $\eta_{4}$ quantifies the influence of $X_{4}$.
+Overall, these results confirm that the proposed given‐data IS framework offers a powerful and cost-effective means to perform detailed distributional local sensitivity analyses of Sobol’ indices.
 5 Conclusion
 In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling.
-We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution different from the reference one.
-In particular, we show that by choosing optimal densities, it is possible to achieve a lower bound on the asymptotic variance than the efficiency bound established in [19].
-In some special cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities.
+We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution that is different from the reference one.
+In particular, we show that by choosing suitable auxiliary densities, the asymptotic variance of efficient estimators can be further reduced, as changing the sampling distribution modifies the efficiency bound itself.
+In some cases, zero variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities.
+In realistic settings where model evaluations are available, one may nevertheless approximate these optimal densities, which naturally leads to significant variance reduction.
 We also show that combining our method with the rank-based estimator [22] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [32].
-Our approach can be naturally extended to other contexts, such as the given-model setting where a computational budget is available to evaluate the black-box model.
-In such cases, we expect to approximate the optimal densities presented in our framework and achieve significant variance reduction.
+The first limitation of our present work is that the optimal densities depend on the code, so additional computational effort is required.
+Another limitation is that we have proved the existence of one optimal distribution for every Sobol’ index, so it may be computationally expensive to estimate all these optimal densities, which increases the numerical effort.
+In a related direction, [15] investigate how several expectations under different target distributions can be estimated from a single Monte Carlo sample, using an adaptive importance sampling scheme with shared control variates.
+This suggests that multi–target designs could also be explored to construct sampling densities efficient for several Sobol’ quantities simultaneously.
 Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
-Appendix A Lemma 1 and other derivations
-Proof of Lemma 1
-We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
-$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
+Acknowledgement
+We gratefully acknowledge our ONERA colleagues — Lokman Bennani, Ghislain Blanchard, Maxime Bouyges, and Patricia Klotz — for providing the ice-accretion dataset.
+We also thank the reviewers for their valuable remarks during the revision process.
+Supplementary Materials
+All simulation codes for the academic experiments (Section 4) are available in the Supplementary Materials and on GitHub:
+$\text{{https://github.com/HaythemBC/Importance-sampling-for-Sobol-indices-estimation}}.$
+Data Availability Statement
+All data used in the academic case studies are generated by simulation and can be fully reproduced using the associated code.
+The real-world dataset of Section 4.3 is confidential and cannot be shared publicly, but can be provided to reviewers upon reasonable request, in accordance with the data-sharing agreement of the contributing institution.
+Appendix A Lemma 3.1 and other derivations
+Proof of Lemma 3.1. We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
+$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
 $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{p}\left[\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
-$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}),$
+$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m(\mathbf{X}_{u}),$
 and
-$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m_{u}^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m_{u}^{2}(\mathbf{X}_{u})]$
-Derivation of the optimal variance under $p_{u}$ in (3.7)
-We start from the expression
-$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$
-Using the last identity we deduce
-$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}=w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u}),\quad\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}=w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u}).$
-For the second term $\mathbb{E}_{q_{\bar{u}}}[Z_{u}^{2}\mid\mathbf{X}_{u}]$, we write
-$\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]$ $=\mathbb{E}_{q}\left[\left(\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y\right)^{2}\middle|\mathbf{X}_{u}\right]$
-$=w_{u}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
-Combining all terms, we obtain:
-$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
-$\quad-3\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
-$=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
-$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
-Computing optimal variance bounds
-Plugging $q^{*}_{\bar{u}\mid u}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{\bar{u}\mid u}))$.
+$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m^{2}(\mathbf{X}_{u})]$
+This completes the proof.
+Derivation of the optimal variance under $p$ (Eq. (12)). From Lemma 3.1, we recall
+$\mathbb{E}_{q}[Z_{u}\mid X_{u}]=\sqrt{w_{u}(X_{u})}\,m(X_{u}),\qquad\bigl(\mathbb{E}_{q}[Z_{u}\mid X_{u}]\bigr)^{2}=w_{u}(X_{u})m^{2}(X_{u}).$
+$\nu_{q}(X_{u}):=\mathbb{E}_{p}[w_{-u\mid u}\,\phi^{2}(X)\mid X_{u}]-m^{2}(X_{u}).$
+We now compute the conditional second moment. Using the definition of $Z_{u}$ and $\phi^{2}(x)$, we obtain
+$\mathbb{E}_{q}[Z_{u}^{2}\mid X_{u}]$ $=\mathbb{E}_{q}\!\left[w_{u}(X_{u})w_{-u\mid u}(X_{-u}\mid X_{u})Y^{2}\mid X_{u}\right]$
+$=w_{u}(X_{u})\,\mathbb{E}_{p}\!\left[w_{-u\mid u}(X_{-u}\mid X_{u})\,\phi^{2}(X)\mid X_{u}\right].$
+Hence
+$\operatorname{Var}_{q}(Z_{u}\mid X_{u})=w_{u}(X_{u})\Big(\mathbb{E}_{p}[w_{-u\mid u}\,\phi^{2}(X)\mid X_{u}]-m^{2}(X_{u})\Big)=w_{u}(X_{u})\,\nu_{q}(X_{u}),$
+Substituting all expressions into Proposition 3.2 yields
+$\sigma^{2}_{\mathrm{opt},u}(q)$ $=\mathbb{E}_{q}\!\left[4\,w_{u}(X_{u})m^{2}(X_{u})\,w_{u}(X_{u})\nu_{q}(X_{u})+w_{u}^{2}(X_{u})m^{4}(X_{u})\right]-\eta_{u}^{2}$
+$=\mathbb{E}_{q}\!\left[w_{u}^{2}(X_{u})\bigl(4m^{2}(X_{u})\nu_{q}(X_{u})+m^{4}(X_{u})\bigr)\right]-\eta_{u}^{2}.$
+Using the usual identity $\mathbb{E}_{q}[w_{u}^{2}g]=\mathbb{E}_{p}[w_{u}g]$, we finally obtain
+$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\!\left[w_{u}(X_{u})\bigl(4m^{2}(X_{u})\nu_{q}(X_{u})+m^{4}(X_{u})\bigr)\right]-\eta_{u}^{2},$
+which is Eq. (12) of the main text.
+Computing optimal variance bounds.
+Plugging $q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{-u\mid u}))$.
 We obtain:
@@
 $=\left(\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\right)^{2}-\eta_{u}^{2}.$
-Derivation of the zero-variance formula
-We start from the general expression:
-$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
-Now Plugging both optimal densities simultaneously:
-$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}^{2}(\mathbf{x}_{u})}$
+Derivation of the zero variance formula We start from the general expression:
+$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
+Now plugging both optimal densities simultaneously.
+$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{-u}^{*}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u}(\mathbf{x}_{-u})\,f(\mathbf{x})}{m^{2}(\mathbf{x}_{u})}$
 $\sigma^{2}_{\mathrm{opt},u}(q)$ $=\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\left(4\,m^{2}(\mathbf{X})\cdot\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}\right]-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
@@
 $=\eta_{u}^{2}-\eta_{u}^{2}=0.$
+Appendix B. Optimal sampling densities for the rank estimator
+In this appendix, we derive the optimal sampling densities for the rank-based estimator of $\eta_{u}=\mathbb{E}[m^{2}(\mathbf{X}_{u})]$.
+Since this estimator is not asymptotically efficient, the optimal variance bound of Section 3 does not apply.
+Nevertheless, the same two-step optimisation over $q_{-u\mid u}$ and $q_{u}$ can be performed, leading to sampling densities specifically adapted to its asymptotic variance.
+Throughout, we use $m(\mathbf{x}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}=\mathbf{x}_{u}]$, $\nu(\mathbf{x}_{u})=\operatorname{Var}(Y\mid\mathbf{X}_{u}=\mathbf{x}_{u})$, and $\phi^{2}(\mathbf{x})=\mathbb{E}[Y^{2}\mid\mathbf{X}=\mathbf{x}]$.
+We first recall the asymptotic variance under the reference distribution:
+$n\,\operatorname{Var}(\widehat{\eta}_{\mathrm{rank},u})=4\,\mathbb{E}\big[m^{2}(\mathbf{X}_{u})\,\nu(\mathbf{X}_{u})\big]+\mathbb{E}\big[\nu^{2}(\mathbf{X}_{u})\big]+\operatorname{Var}\big(m^{2}(\mathbf{X}_{u})\big)+o(1),$
+which can be rewritten as
+$\sigma^{2}_{\mathrm{rank},u}(p)=\mathbb{E}_{p}\big[4\,m^{2}(\mathbf{X}_{u})\,\nu_{p}(\mathbf{X}_{u})+m^{4}(\mathbf{X}_{u})+\nu_{p}^{2}(\mathbf{X}_{u})\big]-\eta_{u}^{2}.$
+Following exactly the derivation of Proposition 3.2, the variance under any density $q$ can be written as
+$\sigma^{2}_{\mathrm{rank},u}(q)=\mathbb{E}_{p}\!\big[w_{u}(\mathbf{X}_{u})\,S_{\mathrm{rank},q}(\mathbf{X}_{u})\big]-\eta_{u}^{2},\qquad\text{where}\qquad S_{\mathrm{rank},q}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{q}(\mathbf{x}_{u})+\nu_{q}^{2}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u}),$
+and
+$\nu_{q}(\mathbf{x}_{u})=\mathbb{E}_{p}\!\left[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}\right]-m^{2}(\mathbf{x}_{u}).$
+As in Section 3, $S_{\mathrm{rank},q}(\mathbf{x}_{u})$ depends on $q_{-u\mid u}$ only through $\nu_{q}(\mathbf{x}_{u})$.
+Hence minimising $S_{\mathrm{rank},q}(\mathbf{x}_{u})$ reduces to minimising $\nu_{q}(\mathbf{x}_{u})$, and the optimal conditional density is exactly the same as in (13):
+$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,\phi(\mathbf{x})}{\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]}.$
+Under this density,
+$\nu_{q^{*}}(\mathbf{x}_{u})=\big(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]\big)^{2}-m^{2}(\mathbf{x}_{u}).$
+Case A and Case B
+As in Section 3, we distinguish the two choices for the conditional component:
+$S_{A}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{p}(\mathbf{x}_{u})+\nu_{p}^{2}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u}),\qquad\text{when }q_{-u\mid u}=p_{-u\mid u},$
+and
+$S_{B}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{q^{*}}(\mathbf{x}_{u})+\nu_{q^{*}}^{2}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u}),\qquad\text{when }q_{-u\mid u}=q^{*}_{-u\mid u}.$
+Compared to Section 3, the only difference is the additional term $\nu^{2}$ in both $S_{A}$ and $S_{B}$.
+For a fixed conditional $q_{-u\mid u}$ (either Case A or Case B), the variance can be written as
+$\sigma^{2}_{\mathrm{rank},u}(q)=\mathbb{E}_{p}\!\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
+where $S$ denotes $S_{A}$ or $S_{B}$ depending on the chosen conditional component. As in Section 3, the optimal marginal is
+$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,\sqrt{S(\mathbf{x}_{u})}}{\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{rank},u}(q_{u}^{*},q_{-u\mid u})=\left(\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\right)^{2}-\eta_{u}^{2}.$
+Resulting optimal density.
+The optimal sampling density for the rank estimator is therefore
+$q^{*}_{\mathrm{rank}}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$
+with the conditional component identical to Section 3, while the marginal depends on $S_{A}$ or $S_{B}$ through the additional term $\nu^{2}$, which encodes the intrinsic sub-efficiency of the rank estimator.

### v3

@@
-Importance sampling for Sobol’ indices estimation
+arXiv:2507.05958v3[math.ST] 01 Mar 2026
 Abstract
@@
 Thanks to the Hoeffding decomposition [27], Sobol’ indices measure the contribution of each group of inputs to the output variance.
-The first-order Sobol’ indices quantify the main effects of each input, while the total-order Sobol’ indices also account for both main and interaction effects.
+The first-order Sobol’ indices quantify the main effect of each input, while the total-order Sobol’ indices also account for interaction effects.
 Hence, they are variance-based measures that capture the contribution of inputs to the overall variability of the model output.
@@
 In particular, for first-order Sobol’ indices, consistency and asymptotic normality have been proven for kernel estimators [12, 11], the Pick–Freeze estimator [28], nearest-neighbor methods [16], and rank-based estimators [22].
-With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, as it requires two evaluations of the model for each realization of $X$, which is often prohibitive.
+With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, since the number of model evaluations required to maintain a given accuracy for all indices increases linearly with the number $k$ of input variables.
 When the context allows it, using this particular design can considerably simplify estimation, but the number of code calls often makes it prohibitive.
@@
 These estimators, however, remain difficult to implement in practice.
-In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalized to attain the efficiency bound in [31] (Proposition 3.3).
+In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalised to attain the efficiency bound in [31] (Proposition 3.3).
 The question of efficiency bounds has been thoroughly discussed in [30].
@@
 In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling.
-More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is the most challenging term to estimate accurately.
-We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution, which makes it possible to characterize optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
+More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators is the most challenging term to estimate accurately.
+We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution,which makes it possible to characterise optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
 To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices.
@@
 $\hat{I}^{\mathrm{IS}}=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{X}^{(i)})\,w(\mathbf{X}^{(i)}).$ (4)
-This estimator achieves zero variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere.
+This estimator achieves zero-variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere.
 This ideal distribution cannot be used in practice because it depends on the unknown quantity to be estimated.
@@
 Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values.
-The next section extends the classical IS framework to characterize the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
+The next section extends the classical IS framework to characterise the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
 3 Main results
@@
 For any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\mathbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x}_{u})$ its marginal density.
-We introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
+We introduce an auxiliary joint density $q(\mathbf{x})$, factorised as
 $q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\,q_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$ (5)
@@
 Let $Z_{u}$ be defined as in Lemma 3.1. Then, for the sampling distribution $q$, the optimal asymptotic variance is given by
-$\boxed{\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.}$ (11)
+$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.$ (11)
 Substituting the expressions of $Z_{u}$, $\sigma^{2}_{\mathrm{opt},u}(q)$ can be equivalently expressed under $p$ as
@@
 In the second step, the resulting objective is minimised with respect to $q_{u}$.
-To characterize the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations.
+To characterise the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations.
 Following the classical Lagrangian minimisation framework, one can formally derive the stationary condition satisfied by $q^{*}$ under the normalisation constraint.
 In general, verifying that this stationary point corresponds to a minimum would require checking second-order optimality conditions.
-However, as mentioned by Owen [34, End Notes, p.
+However, as mentioned by Owen [35, Chapter 9, End Notes (“Calculus of variations”), p.
 38], the Cauchy–Schwarz inequality is sufficient to establish global minimality in this setting.
@@
 Joint optimisation in Case B therefore achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
-3.4 Zero variance optimal distribution
+3.4 zero-variance optimal distribution
 In the idealised deterministic case, where all inputs are controllable and no external noise variable $\mathbf{W}$ is present ($Y=f(\mathbf{X})$), it is theoretically possible to achieve zero variance in estimating $\eta_{u}$.
@@
 $\sigma^{2}_{\mathrm{opt},u}(q^{*})\leq\sigma^{2}_{\mathrm{opt},u}(p).$
-In general, this distribution does not yield a zero variance estimator, since a direct plug-in calculation shows
+For instance, this distribution does not yield a zero-variance estimator, since a direct plug-in calculation shows
 $\sigma^{2}_{\mathrm{opt},u}(q^{*})=\left(\mathbb{E}_{p}\!\left[\sqrt{S_{C}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}>0.$
 Remark.
-When the model $f(\mathbf{x})$ may take both positive and negative values, the theoretical zero variance density derived above is no longer a valid probability distribution.
-A practical alternative consists in separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution.
-This idea, sometimes referred to as positivisation in the importance sampling literature [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
+When the model $f(\mathbf{x})$ takes both positive and negative values, the associated optimal variance is strictly positive, so the zero-variance property cannot be achieved.
+If one nevertheless wishes to construct a zero-variance estimator, a classical alternative consists of separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution.
+This idea, referred to as positivisation [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
 We decompose $f$ into its positive and negative parts:
@@
 Two auxiliary densities $q_{+}$ and $q_{-}$ are then defined to be positive wherever $pf_{+}$ and $pf_{-}$ are positive, respectively.
-Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, preserving unbiasedness while reducing variance in practice.
+Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, yielding an unbiased estimator.
+In the ideal case where each component is sampled according to its fully adapted optimal density, the resulting estimator achieves zero variance.
 If $f(\mathbf{x})\geq 0$ for all $\mathbf{x}$, then $\mathbb{E}_{p}[|f(\mathbf{X})|\mid\mathbf{X}_{u}]=m(\mathbf{X}_{u})$ and $S_{C}(\mathbf{x}_{u})=m^{4}(\mathbf{x}_{u})$, so that
@@
 The resulting fully adapted sampling density is
-$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
-In theory, the density $q^{*}(\mathbf{x})$ yields a zero variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it.
+$q_{u}^{*}(\mathbf{x})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
+In theory, the density $q^{*}(\mathbf{x})$ yields a zero-variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it.
 Adaptive procedures can be employed, including nonparametric methods [50] and parametric ones such as the cross-entropy method [39, 14].
@@
 This function is a standard benchmark in sensitivity analysis due to its nonlinear, non-monotonic structure and the availability of analytical Sobol’ indices.
-We consider the 3-dimensional case $(k=3)$ with $a_{i}=i$ and focus on the second-order index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$.
-For this part, one could in principle rely on kernel estimators [18, 11, 12], but since all quantities are analytically tractable, we rely on their closed forms and focus on the impact of optimal importance densities on variance reduction.
+We consider the 3-dimensional case $(k=3)$ with $a_{i}=i$ and focus on the second-order index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$.For this part, one could in principle rely on kernel estimators [18, 11, 12], but since all quantities are analytically tractable, we rely on their closed forms and focus on the impact of optimal importance densities on variance reduction.
 The third input $X_{3}$ is treated as an uncontrolled variable $W$.
@@
 In particular, we show that by choosing suitable auxiliary densities, the asymptotic variance of efficient estimators can be further reduced, as changing the sampling distribution modifies the efficiency bound itself.
-In some cases, zero variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities.
+In some cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities.
 In realistic settings where model evaluations are available, one may nevertheless approximate these optimal densities, which naturally leads to significant variance reduction.
@@
 $=\left(\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\right)^{2}-\eta_{u}^{2}.$
-Derivation of the zero variance formula We start from the general expression:
+Derivation of the zero-variance formula We start from the general expression:
 $\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
@@
 The optimal sampling density for the rank estimator is therefore
-$q^{*}_{\mathrm{rank}}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$
+$q_{u}^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$
 with the conditional component identical to Section 3, while the marginal depends on $S_{A}$ or $S_{B}$ through the additional term $\nu^{2}$, which encodes the intrinsic sub-efficiency of the rank estimator.
+Deterministic case ($Y=f(\mathbf{X})$).
+Assume that $f\geq 0$ and that there is no stochastic noise, i.e.
+$Y=f(\mathbf{X})$.
+Then $\phi^{2}(\mathbf{x})=\mathbb{E}[Y^{2}\mid\mathbf{X}=\mathbf{x}]=f^{2}(\mathbf{x})$ and thus $\phi(\mathbf{x})=f(\mathbf{x})$.
+In particular,
+$m(\mathbf{x}_{u})=\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}],\qquad\nu_{p}(\mathbf{x}_{u})=\operatorname{Var}_{p}(f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}).$
+The optimal conditional density becomes
+$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]}=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{m(\mathbf{x}_{u})}.$
+Moreover, under $q^{*}_{-u\mid u}$ we have
+$\nu_{q^{*}}(\mathbf{x}_{u})=\big(\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]\big)^{2}-m^{2}(\mathbf{x}_{u})=0,$
+so that $S_{B}(\mathbf{x}_{u})=m^{4}(\mathbf{x}_{u})$ and hence
+$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}.$
+Therefore, the resulting optimal sampling density is
+$q_{u}^{*}(\mathbf{x})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}}.$